OpenStudy (prepfortears):

Will medal and fan! What is the value of an annuity due at the end of 15 years of quarterly deposits of $2,000.00 with terms of 8 percent compounded quarterly?

OpenStudy (prepfortears):

Can someone please help with this?

OpenStudy (prepfortears):

I really just need help with how to deal with the 8% compounded quarterly

OpenStudy (amistre64):

annuity due, is that the one at the begining or end of the month?

OpenStudy (amistre64):

i spose ordinary is end, so annuity due starts with a payment

OpenStudy (amistre64):

if we start with a balance of the first payment, and then add on payments for each compounding period, the formula i usually use turns to this \[B_n=Pk^n+P\frac{k^n-1}{k-1}\]

OpenStudy (amistre64):

i dont know what formula you use tho ....

OpenStudy (prepfortears):

i've never done a question like this one before

OpenStudy (amistre64):

what types have you done?

OpenStudy (amistre64):

how do you find the value of an ordinary annuity?

OpenStudy (prepfortears):

i've never worked with annuities before

OpenStudy (amistre64):

well, what have you worked with?

OpenStudy (amistre64):

Have you worked with loans yet?

OpenStudy (amistre64):

monthly payments?

OpenStudy (prepfortears):

nope, i'm in a class and the very first thing is annuities

OpenStudy (amistre64):

can you define what an annuity does? how would you explain it to someone?

OpenStudy (prepfortears):

an amount of money payed out at intervals?

OpenStudy (amistre64):

yeah, so payments

OpenStudy (amistre64):

lets say we pay ourself, we put money into an account that compounds the interest, quarterly and each quarter we add more money .. if we do this for 15 years, how much money do we have saved up in the account? this is the goal we are looking at

OpenStudy (amistre64):

i start with amount P, and it compounds for 15*4 periods i then add another P, that compounds for 15*4-1 periods i then add another P, that compounds for 15*4-2 periods i then add another P, that compounds for .... i then add another P, that compounds for 0 periods, i then look at the balance in my account

OpenStudy (amistre64):

the only difference between an annuity due (starting with a payment) and an ordinary annuity (starting at 0) is that first payment that compounds for the full 15 years

OpenStudy (amistre64):

but I spose you dont know how to add all that up do you

OpenStudy (prepfortears):

no not really sorry

OpenStudy (amistre64):

right, and thats why i provided my own formula, it puts the addition of all the payments into a nice neat format. all i need from you is to tell me what it means to compound once a quarter, what is out interest adjustment?

OpenStudy (amistre64):

what is **our interest adjustment

OpenStudy (amistre64):

8% is a yearly interest amount, we dont compound in a year, but less then a year ... we compound 4 times a year. 8/4 = 2% each time

OpenStudy (amistre64):

1.02 is out compounding rate for each period, and we have 15 years of 4 times a year = 60 periods in total

OpenStudy (amistre64):

do you agree or disagree? any questions?

OpenStudy (prepfortears):

im good so far

OpenStudy (amistre64):

so lets say k=1.02 P = 2000 B0 = P B1 = Pk + P , money compounds and we add a payment B2 = Pk^2 + Pk + P , money compounds and we add a payment B3 = Pk^3 + Pk^2 + Pk + P , money compounds and we add a payment this goes on and on and is a geometric series in k, which there is a nice formula for fining the sum of \[B_n=Pk^n+P\underbrace{\frac{k^n-1}{k-1}}_{geo~series~sum}\]

OpenStudy (amistre64):

i leave the first term where its at in my formula becasue the generally represents a starting balance that is different from P.

OpenStudy (amistre64):

the ordinary annuity starts at a balance of 0, so the first term would be 0 in that case.

OpenStudy (amistre64):

so, what is: 2000(1.02)^(60) + 2000((1.02)^(60)-1)/(.02)

OpenStudy (prepfortears):

234,665.14

OpenStudy (amistre64):

thats what i get, so that is how much the annuity is worth, what is its present value?

OpenStudy (amistre64):

FV = PV(1.02)^(60) if we were to sell it today to an investor, they would pay us the present value of it or less for profit

OpenStudy (amistre64):

234665/(1.02)^(60) = 71522 as a lump sum today for your annuity would have them break even. Its the amount of money they would have to put in the bank earning the same interest for the same time period

OpenStudy (amistre64):

but my memory could be faded on that so dont worry to much about it

OpenStudy (prepfortears):

ok i think i have the hang of this i have another similiar question, would you be able to check it as I work through it to make sure im doing it right?